OpenStudy (anonymous):

A triangular garden is 20 ft taller than it is wide. The area is 750ft^2 . Find the height and the base of the garden.

OpenStudy (anonymous):

First find the formula for an area of a triangle. What is that?

OpenStudy (anonymous):

1/2*b*h=A b=h-20 1/2*h^2-10h=A A=750 (1/2)h^2-10h-750=0 quadratic form will give you your answer.

OpenStudy (anonymous):

1/2bh=A

OpenStudy (anonymous):

1/2b(20+b)=750

OpenStudy (anonymous):

Yes now FOIL that problem out and bring the 750 to the left side so you can use the quadratic equation to solve.

OpenStudy (anonymous):

this is where im stuck because 1/2 and b are together how do i foil with that?

OpenStudy (anonymous):

FOIL it as .5b, think of it as a whole not two different numbers.

OpenStudy (anonymous):

distribute. not foil. .5b*20+.5b*b

OpenStudy (anonymous):

Eh, Foil it without the outer....I still think of it as foil. Distribute isbetter termanology tho

OpenStudy (anonymous):

this right? .(1/2)b*20+(1/2)b*b

OpenStudy (anonymous):

Yes so you will finish with .5b^2 + 10b - 750 = 0...then quadratic equation and you got yourself the answer.

OpenStudy (anonymous):

how do i find the height and base from that?

OpenStudy (anonymous):

Find you base and then plug your base back into your equation and get your height.

OpenStudy (anonymous):

b is the base. h is the height. once you find one you have the equation to find the other.

OpenStudy (anonymous):

but i dont understand how to solve for b in that type of equation...

OpenStudy (anonymous):

quadratic formula. google it.

OpenStudy (anonymous):

Honestly google is your best bet. All you have to do is, plug in your numbers compared to show the quadratic equation needs them and it will show you how to find an answer.

OpenStudy (anonymous):

is it b=-50 and b=30? thats what i keep getting...

OpenStudy (anonymous):

I haven't done the math myself, but seems like reasonable answers, no think can a base length be negative? That will tell you which is the correct number to use to plug back in for h

OpenStudy (anonymous):

so id plug in 30 as b and solve for h that way?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

thanks...